OpenStudy (anonymous):

HELP....!!! WHAT IS THE SLOPE OF THE LINE REPRESENTED BY THE EQUATION BELOW y=-2/3x+3

OpenStudy (eric_d):

y=mX+c

OpenStudy (eric_d):

m=slope

OpenStudy (anonymous):

WHATS THE ANSWER

OpenStudy (eric_d):

-2/3

OpenStudy (anonymous):

your equation in is y=mx+b form where the m is the slope, what is 'm' in your equation?

OpenStudy (tester97):

First stop requesting the answer and let us show you how to get the answer by due process. Secondly use what @eric_d said; y=mx+c but I have seen y=mx+b. Anyways, If your equation is y=-2/3x+3 then the slope is just -2/3

OpenStudy (eric_d):